### v2

# Are LLMs Good Literature Review Writers? Evaluating the Literature Review Writing Ability of Large Language Models

Abstract
The literature review is a crucial form of academic writing that involves complex processes of literature collection, organization, and summarization. The emergence of large language models (LLMs) has introduced promising tools to automate these processes. However, their actual capabilities in writing comprehensive literature reviews remain underexplored, such as whether they can generate accurate and reliable references. To address this gap, we propose a framework to assess the literature review writing ability of LLMs automatically. We evaluate the performance of LLMs across three tasks: generating references, writing abstracts, and writing literature reviews. We employ external tools for a multidimensional evaluation, which includes assessing hallucination rates in references, semantic coverage, and factual consistency with human-written context. By analyzing the experimental results, we find that, despite advancements, even the most sophisticated models still cannot avoid generating hallucinated references. Additionally, different models exhibit varying performance in literature review writing across different disciplines.
1 Introduction
The literature review is a form of academic writing designed to summarize, analyze, and evaluate existing research findings, theoretical perspectives, and methodologies. However, writing a literature review often requires extensive reading and summarizing of relevant literature. Literature reviews in popular fields may require tens or even hundreds of references, which is a complex and labor-intensive task. In recent years, deep learning models have been developed to assist in the automated writing of literature reviews Kontonatsios et al. (2020); Aliyu et al. (2018). The advent of large language models (LLMs) has brought promising advances in automating certain stages of literature review writing. With their powerful generative capabilities, enhanced by training on vast amounts of academic text, LLMs have emerged as a formidable tool for writing literature reviews Wang et al. (2024).
However, how well LLMs perform in writing literature reviews remains unclear. To the best of our knowledge, no comprehensive studies have yet evaluated the literature review writing capabilities of LLMs. Literature review writing requires, on one hand, deep familiarity with the studies in a specific field, and on the other hand, the ability to summarize the core contributions of that literature. While techniques such as Retrieval-Augmented Generation (RAG) can enhance LLMs’ domain-specific knowledge, most researchers use LLMs without RAG such as ChatGPT for literature review writing. Therefore, it is necessary to evaluate the literature review writing capabilities of these native LLMs.
Therefore, in this paper, we propose a framework for automatically assessing literature review writing ability in LLMs. The evaluation of LLMs’ ability to write literature reviews remains an open question. In this study, we use human-written literature reviews as the gold standard and employ external tools for a comprehensive evaluation. Specifically, we first collect human-written literature reviews to create the dataset. Then, we ask LLMs to complete three tasks: generating references, writing an abstract, and writing a literature review based on a given topic. The generated results are then evaluated from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context.
Our contribution can be summarized as follows.
  • First, we propose a framework for automatically evaluating the literature review writing ability of LLMs, without requiring any human involvement. This framework encompasses multiple stages, including dataset construction, LLMs gnenrated data collection, and evaluation.
  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
  • Then, we evaluate the generated results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
  • Finally, we assess four LLMs using the proposed framework. By analyzing the experimental results, we find that hallucinated references remain an unavoidable issue for current LLMs. Additionally, the performance of LLMs in writing literature reviews varies across different disciplines.
2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024) proposed a literature survey writing process for LLMs, called AutoSurvey. By RAG retrieving the latest relevant paper incorporates real-time knowledge. In evaluation, Muti-LLM-as-judge is used to evaluate citation quality and content quality.  Wang_Hu_Wang_Yan_Sheng_He_2024 compared the use of LLMs across different academic tasks and found that researchers are more likely to use LLMs for literature review (LR) tasks than for paper understanding (PU) tasks.
Additionally, more attention has been paid to the hallucination evaluation of reference citations. According to whether the output contradicts the prompt or not,  OpenAI et al. (2023) categorized hallucination phenomenon as open domain and closed-domain hallucinations. Closed-domain hallucination assessments require only comparisons with input contexts, whereas open-domain hallucination assessments require more extensive research Bubeck et al. (2023), such as human evaluation Ji et al. (2023).
Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generated 11 systematic reviews under the shoulder rotator cuff pathology topic, then to evaluate the hallucination rates of reference. They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4. For example,  Agrawal et al. (2024) chose 200 computer science topics to evaluate LLMs’ hallucination rate. They first asked LLMs to generate 5 reference titles, then use BING Search API to label the reference automatically as true or not. The LLMs were then examined to see if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?).
Athaluri et al. (2023) assessed the hallucinations of LLMs in research proposal writing. They used ChatGPT to generate 50 research proposals based on the given 50 topics. ChatGPT generated 178 references in these proposals and then they employed five researchers to manually check these references and DOIs on Scopus, Google and PubMed search engines and found that among the references 109 references had valid DOIs. score  Aljamaan et al. (2024) given 5 medical topics, then used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generated 10 relevant references for each topic, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication data, link to the paper, and relevance of the reference to the keyword prompt. Then, they proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date. The existing studies on evaluating citation hallucinations in literature reviews are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance. Moreover, these evaluations often involve manual human intervention.
3 Methodology
In this section, we present the framework for evaluating LLMs’ literature survey writing ability. The framework as shown in Figure 1, consists of three main stages: first, dataset construction; second, task design for evaluation; and third, assessment of the generated results.
3.1 Dataset Construction
As illustrated in Figure 1, we first collect literature survey data from the Annual Reviews website ( https://www.annualreviews.org/). Annual Reviews is an independent, nonprofit scholarly publishing company, which publishes 51 journals of review articles. The articles on the website are peer-reviewed, so they are of high quality and can be used as experimental materials. We crawl all articles published in 2023, including their title, keywords, abstracts, contents, and references. Then clean them as the experimental dataset. Assessing the ability to write literature reviews is a challenging task, as evaluating the quality of content is inherently complex. In this paper, we use human-written reviews as the gold standard, which simplifies the evaluation process to some extent.
Then, the dataset $D$ is the article set from $K$ journals, $D=\{j_{0},...,j_{k},...,j_{K}\}$, the number of journals is 51, so $K=50$. For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles in $j_{k}$, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
3.2 Task Design
We design three tasks as follows.
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract according to the research topic provided in the title and the keywords. The length of the generated abstract is required to be the same as the length of the original abstract of the article. In this task, we evaluate whether LLMs can write an abstract based on the given topic.
  • Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a literature review according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies (with a total of $N$ citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
Three task prompts are shown in Appendix Table LABEL:at1.
3.3 Evaluation Metrics
Based on the type of generated text, we divide the evaluation of the model’s results into two parts: first, the hallucination rate of the references generated by LLMs, and second, a comparison of the generated context with human-written results, including two dimensions: factual consistency and semantic coverage.
Figure 1: Illustration of the proposed framework.
Reference hallucination evaluation metrics. Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate accurate references. In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM.
For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task1 and Task3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the article title, authors, journal, year, volume, first page, and last page. Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
Next, we describe how to obtain $\{e_{d}\mid d=0,1,\dots,7\}$. First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines. Then, we obtain two candidate sets $C_{t}$ and $C_{a}$ respectively. Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $C_{t}$ and $C_{a}$. For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$. Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $c_{n}^{g}$ of $r_{n}^{g}$.
Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $c_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1). On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article. Here its title $T$ is correct or not, we give LLM a certain margin of error, as long as the title has a match rate of 80%, it will be judged as accurate, so we need another piece of information that is correct to ensure that the article is reliable. On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article. therefore, when the other three elements in $r_{n}^{g}$ are correct, we also consider it as a reliable article.
$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}$ (1)
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, for each journal $j_{k}$, we can obtain the accuracy of generated references $Acc_{ref.}(j_{k})$ based on Eq 3. Subsequently, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(j_{k})$ across journals, as shown in Eq 4. Finally, we defined the reference accuracy metric $Acc$ as Eq 4. A higher accuracy metric indicates a lower hallucination rate.
$Acc_{ref.}(p_{i})=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})$ (2)
$Acc_{ref.}(j_{k})=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
$Acc.=\frac{1}{K}\sum_{k=0}^{K}{Acc_{ref.}(j_{k})}$ (4)
On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness. In  Agrawal et al. (2024)’s work, they determined by ground-truth labels assigned using the Bing search API. Encouraged by this work, we also calculate the title search scores of each journal and each LLM, formulated as follows.
$S_{t}=1/(NMK)\sum_{K}\sum_{M}\sum_{N}s_{p_{i}}$ (5)
$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from Semantic Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (6)
where $s$ is the number of correct generated reference titles for $p_{i}$. We search the generated title from an external academic search engine to estimate whether the title can be matched.
Context evaluation metrics. In this paper, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects. The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al. (2023). Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text. For example, we calculate the NLI score $Entail_{i}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
$Entail_{i}=\theta_{\text{NLI}}(a_{i},a_{i}^{g})=\begin{cases}1&\text{if}a_{i}^{g}\text{ entails }a_{i}\text{,}\\ 0&\text{otherwise}\end{cases}$ (7)
where $\theta_{NLI}$ denotes the NLI model. Finally, we obtain the NLI score $Entail$ for each model according to Eq 8.
$Entail=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}{Entail_{i}}$ (8)
Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context. Specifically, for Task 2, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation. For Task 3, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
KPR, first proposed by  Qi_Xu_Guo_Wang_Zhang_Xu_2024, is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents. Since the human-written literature review is lengthy and difficult to compare directly, we adapt the KPR metric by first using GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculating the coverage of these key points by the model-generated literature review as Eq. 9.
$KPR=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (9)
where $c_{i}^{g}$ denotes the literature review generated by LLMs. Finally, we also concatenate key points and compute the ROUGE metric between the key point and $c_{i}^{g}$.
4 Experiments
4.1 Experimental Settings
Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1105 articles, the dataset distribution across journals as shown in Appendix B Figure 4.
LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3 and Llama-3.2-3B-Instruct.
In task 1 and task 3, we set $N$ as 10, each model generates 10 references. For the generated reference evaluation, we use Semantic Scholar as the external database. Recently, LLMs-as-judges has become more common Chen et al. (2024); Zheng et al. (2023); Shangyu et al. (2024). So, for Task 2, we employ TRUE Honovich et al. (2022), along with GPT-4o as NLI models for evaluating factual consistency in context; To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings. For task 3, we employ GPT-4o as the NLI model in Eq 9, and set $q$ as 10.
4.2 Main Results
We present the results for the three tasks in Table 12, and 3. The performance of different models on each task is analyzed as follows.
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5 achieves the highest accuracy and title search rate, whereas Llama-3.2 exhibits the lowest performance on both metrics for Task 1. Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs. Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2. The graph shows that the majority of the correctly cited references are concentrated between 2010 and 2020.
Results for Task 2. We use various metrics to evaluate model performance. As shown in Table 2, Claude-3.5 consistently achieves the best overall performance across nearly all metrics. The average semantic similarity between the abstracts generated by Claude-3.5 and human-written abstracts is 81.17%. Furthermore, Claude-3.5 demonstrates the highest factual consistency, with a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.77%. Additionally, we observe that Llama-3.2 achieves the highest score on the ROUGE-L metric, but does not show a clear advantage on other metrics. Given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
Results for Task 3. As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3. Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text. In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly. This suggests that the generated references and the literature review text impose constraints on each other.
On the other hand, we find that Claude-3.5 achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5 recalls the most claims from the human-written literature review. While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
  Models Task 1
$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
Qwen-2.5-72B 21.80 12.25 17.58
Llama-3.2-3B 16.62 3.45 6.95
DeepSeek-V3 56.04 46.33 50.66
GPT-4o 32.07 21.65 24.65
Claude-3.5-Sonnet 64.82 51.59 55.77
Table 1: The experimental results of the four LLMs in Task 1. “$S_{t}$” refers to the title search rate as defined in Eq 5, while “Acc.” represents the accuracy defined in Eq 4. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task 2
[…]
Qwen-2.5-72B 80.22 69.52 95.02 40.61 8.78 20.12
Llama-3.2-3B 79.28 62.39 92.14 40.35 8.96 20.52
DeepSeek-V3 80.96 78.55 96.84 41.13 8.98 20.33
GPT-4o 80.96 77.91 96.50 40.70 8.56 19.86
Claude-3.5-Sonnet 81.17 78.90 96.77 41.13 8.99 20.00
Table 2: Compare the performance of four LLMs on Task 2.
  Models Task 3
References Literature Review
[…]
Qwen-2.5-72B 40.02 28.91 33.64 38.82 29.95 9.01 15.14
Llama-3.2-3B 21.78 4.86 7.28 29.07 28.07 7.77 15.46
DeepSeek-V3 62.29 52.81 55.38 56.02 35.65 10.40 17.46
GPT-4o 60.05 50.62 54.16 59.18 30.78 9.72 15.54
Claude-3.5-Sonnet 66.43 59.06 63.06 62.32 28.59 8.90 14.41
Table 3: The experimental results of the four LLMs in Task 3. “$S_{t}$” refers to the title retrieval rate as defined in Eq 5.“Acc.” represents the accuracy defined in Eq 5. “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task1 Task3
Title Acc. Overlap Overlap(first author) Title Acc. Overlap Overlap(first author)
Qwen-2.5-72B 7.20 12.60 13.27 12.15 17.36 18.31
Llama-3.2-3B 5.97 8.48 8.67 8.01 8.28 8.44
DeepSeek-V3 14.28 19.72 20.50 20.44 26.79 27.30
GPT-4o 8.66 18.76 19.50 18.31 27.88 28.86
Claude-3.5-Sonnet 16.06 24.34 25.21 21.20 31.90 33.25
Table 4: Overlap rate of citations between LLM-generated and human-written references in Task 1 and Task 3. “Title Acc.” refers to the title overlap between LLM-generated and human-written references, while “overlap” represents the article overlap between LLM-generated and human-written references. “first author” refers to when evaluating the overlap of references, the author dimension only comparing the first author.
Figure 2: Distribution of LLM-generated true references over years.
4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5 demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5, GPT-4, and Qwen-2.5 follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. Accuracy for journal name, page, and author is also relatively high. However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions. In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions. However, overall, Llama-3.2 does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5, DeepSeek-V3, GPT-4o, and Qwen-2.5, with particularly obvious gains in the author dimension. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension. Notably, the accuracy of DeepSeek-V3 and GPT-4 in certain dimensions approaches or even exceeds that of Claude-3.5. However, the performance of LLaMA-3.2 remains suboptimal.
4.4 Cross-Disciplinary Analysis
In this section, we compare the performance of LLMs across different disciplines. First, based on Dewey’s Decimal Classification, we categorize 51 journals into five disciplines: Biology, Chemistry, Mathematics, Physics, Social Science, and Technology. After categorization, there are 460 articles in the Biology category, 90 in Chemistry, 50 in Mathematics, 113 in Physics, 299 in Social Science, and 94 in Technology.
We then present bar charts in Figures  LABEL:f4, which illustrate the performance of different models across various tasks and disciplines.
First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline. Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5 (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5 (p<.0001), Llama-3.2 (p=0.065). These results indicate significant differences in performance across disciplines for each model except Llama-3.2.
Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5 (p<.001), Llama-3.2 (p<.0001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
Additionally, we examine the accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c. It is evident that the accuracy of Claude-3.5, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5 and LLaMA-3.2 across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5(p<.0001), Llama-3.2(p<.001), indicating significant differences in each models’ performance across the various disciplines.
Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d. The results from the figure indicate that the differences between models—Claude-3.5, DeepSeek-V3, GPT-4, and Qwen-2.5—are not significant across various disciplines, a finding that is also supported by statistical tests. Specifically, the p-values are as follows: Claude-3.5 (p=0.46), DeepSeek-V3 (p=0.23), GPT-4 (p=0.18), Qwen-2.5 (p=0.10), and Llama-3.2 (p<0.001).
4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5 are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. Next, we provide a more detailed comparison of the references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5 and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
5 Conclusion
In this paper, we present a framework to assess the literature review writing abilities of LLMs. This framework includes three tasks designed to evaluate LLMs’ literature review writing capabilities. The generated outputs are then evaluated from multiple dimensions using various tools, such as Semantic Scholar and NLI models, focusing on aspects like hallucination rate, semantic coverage, and factual consistency compared to human-written texts. Finally, we analyze the performance of LLMs in writing literature reviews from the perspective of different academic disciplines.
We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms DeepSeek-V3, GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references. This advantage is likely influenced by the training data of each model. Additionally, we observed that each model has different strengths across disciplines. Overall, for the reference generation task, nearly all models perform better in Mathematics, while their performance is weaker in Chemistry and Technology. However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.
When comparing the references generated by the models in Task 1 and Task 3, we find that in Task 3, nearly all models generate more accurate references. This suggests that LLMs cite references during the writing process, which improves the authenticity of the references. Moreover, the inclusion of the first author’s name in the generated context also enhances the accuracy of the author dimension.
In the future, we will deploy this evaluation framework on Hugging Face to enable real-time assessments of newly released LLMs.
6 Limitations
In this paper, we evaluate the ability of LLMs to write literature reviews. However, instead of assessing from conventional perspectives, such as the fluency of the generated context or topic coverage, we compare the model-generated results with human-written ones. Additionally, when processing the LLM-generated results, we often need to handle abbreviations of author names and journal titles. While we have made efforts to address these details thoroughly, some discrepancies may still exist.
Appendix A Prompts for Tasks
Table 5: Prompts for tasks.
Prompt Content
Prompt 1 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to find the 10 studies that are most relevant to the research topic provided in the "Title" and the "Keywords" below.
Please cite the studies according to the following JSON format. There is no need to provide any explanation before or after the JSON output. Ensure that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.". { "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
Title: title
Keywords: keywords
Prompt 2 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write an abstract according to the research topic provided in the "Title" and the "Keywords" below. Please write the abstract for about xx words, according to the JSON format as follows. There is no need to provide any explanation before or after the JSON output. {"Abstract": ""}
Title: title
Keywords: keywords
Prompt 3 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write a literature review according to the research topic provided in the "Title", “Abstract” and "Keywords" below. The literature review should be about 1000 words long. I would like you to back up claims by citing previous studies (with a total of 10 citations in the literature review). The output should be in JSON format as follows: { "Literature Review": "xxx", "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
The "Literature Review" field should be about 1000 words. The "References" field is a list of 10 references, and ensures that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.".
Title: title
Keywords: keywords
Abstract: abstract
Appendix B Data Distribution
Figure 4: Statistics of dataset.