### v3

# Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition

Abstract
Large language models (LLMs) have emerged as a potential solution to automate the complex processes involved in writing literature reviews, such as literature collection, organization, and summarization. However, it is yet unclear how good LLMs are at automating comprehensive and reliable literature reviews. This study introduces a framework to automatically evaluate the performance of LLMs in three key tasks of literature writing: reference generation, literature summary, and literature review composition. We introduce multidimensional evaluation metrics that assess the hallucination rates in generated references and measure the semantic coverage and factual consistency of the literature summaries and compositions against human-written counterparts. The experimental results reveal that even the most advanced models still generate hallucinated references, despite recent progress. Moreover, we observe that the performance of different models varies across disciplines when it comes to writing literature reviews. These findings highlight the need for further research and development to improve the reliability of LLMs in automating academic literature reviews.
1 Introduction
The literature review is a critical component of academic writing that aims to synthesize, critique, and assess the current state of knowledge in a particular field. It involves a comprehensive examination of published research articles, theoretical frameworks, and research methodologies related to a specific topic. Conducting a thorough literature review often necessitates extensive reading and summarizing of pertinent literature, which can be a complex and time-consuming process, especially in well-established fields where the number of relevant references can range from dozens to hundreds. To alleviate this burden, researchers have recently turned to advanced deep learning models as a potential tool to aid in the automated generation of literature reviews. Kontonatsios et al. (2020); Aliyu et al. (2018). The emergence of large language models (LLMs) has introduced a promising avenue for automating various aspects of the literature review writing process. LLMs, trained on extensive collections of academic texts, have developed impressive generative capabilities that make them a potent tool for assisting researchers in crafting literature reviews. By leveraging the knowledge acquired from vast amounts of scholarly literature, these models can help researchers quickly identify relevant sources, summarize key findings, and generate coherent syntheses of the existing body of work Wang et al. (2024b); Agarwal et al. (2024).
The process of writing a literature review requires a deep understanding of the studies within a particular field, as well as the ability to summarize the core contributions of the relevant literature. While techniques such as Retrieval-Augmented Generation (RAG) can enhance the domain-specific knowledge of LLMs, most researchers rely on vanilla LLMs, such as ChatGPT, for literature review writing without the use of RAG Wang et al. (2024a). Consequently, it is crucial to evaluate the performance of these native LLMs in the context of literature review writing to determine their effectiveness and limitations.
Therefore, in this paper, we propose a framework for automatically assessing the literature review writing ability of LLMs, using human-written literature reviews as the gold standard and designing metrics for a comprehensive evaluation. We first collect a dataset of human-written literature reviews to serve as a benchmark for evaluating the performance of LLMs. We then ask LLMs to complete three tasks based on the collected dataset: generating references, writing an abstract, and writing a complete literature review based on a given topic. Finally, we evaluate the generated results from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context. By assessing the performance of LLMs across these tasks and evaluating their output using our proposed metrics, we aim to provide a comprehensive understanding of their capabilities and limitations in writing literature reviews.
Our contribution can be summarized as follows.
  • First, we propose a framework for automatically evaluating the literature review writing ability of LLMs, without requiring any human involvement. This framework encompasses multiple stages, including the compilation of a literature review dataset construction, the collection of LLM-generated output, and the evaluation of LLM performance.
  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks for accessing LLMs in literature writing: reference generation, literature summary, and literature composition on a given topic.
  • Then, we evaluate the generated results of LLMs from multiple perspectives, including the hallucination rate in generated references, factual consistency, and semantic coverage compared to human-written content.
  • Finally, we assess five LLMs using the proposed framework. By analyzing the experimental results, we find that hallucinated references remain a prevalent issue for current LLMs. Furthermore, the performance of LLMs in writing literature reviews varies across different disciplines.
2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024b) proposed a literature survey writing process for LLMs, called AutoSurvey, which incorporates real-time knowledge by retrieving the latest relevant papers using a retrieval-augmented generation (RAG) approach. Similarly,  Agarwal et al. (2024) investigated the zero-shot capabilities of recent LLMs in generating literature reviews from abstracts. They employed a two-step approach: first retrieving relevant papers via an external knowledge base, then outlining and executing a structured plan to generate the review. The authors focused on the LLMs’ ability to effectively summarize the literature.
Additionally, for evaluating the literature review writing capabilities of LLMs, more attention has been paid to assessing hallucinations in reference citations. For example, Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generate 11 systematic reviews on shoulder rotator cuff pathology and evaluated the hallucination rates of the references. They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4. Similarly, Agrawal et al. (2024) chose 200 computer science topics to evaluate LLMs’ hallucination rates. They first asked LLMs to generate 5 reference titles, then used the BING Search API to automatically label the references as true or not. The LLMs were then examined to determine if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?). Athaluri et al. (2023) assessed the hallucinations of LLMs in research proposal writing. They used ChatGPT to generate 50 research proposals based on 50 given topics. ChatGPT generated 178 references in these proposals, and then they employed five researchers to manually check these references and DOIs on Scopus, Google, and PubMed search engines. They found that among the references, 109 had valid DOIs. Moreover, Aljamaan et al. (2024) provided 5 medical topics and used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication date, link to the paper, and relevance of the reference to the keyword prompt. They then proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date. While these studies are very useful attempts at evaluating citation hallucinations in literature reviews, they are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance. Moreover, these evaluations often involve manual human intervention.
3 Methodology
In this section, we propose a framework for evaluating LLMs’ literature survey writing ability. The framework, as shown in Figure 1, consists of three main stages: dataset construction and task design for evaluation, collection LLM-generated output, and assessment of the generated output.
3.1 Dataset Construction
Assessing the ability to write literature reviews is a challenging task, as evaluating the quality of content is inherently complex. In this paper, we use human-written reviews as the gold standard, which simplifies the evaluation process to some extent. As illustrated in Figure 1, we first collect publicly available information of literature reviews (i.e., the title, authors, abstract, and keywords) from the Annual Reviews website (https://www.annualreviews.org/). Annual Reviews, an independent nonprofit publisher, produces 51 review journals spanning various scientific disciplines. Invited experts write comprehensive, authoritative reviews that synthesize and summarize the most significant primary research literature in their field, providing a valuable resource for researchers to stay current with the latest developments. We crawl all articles published in 2023, including their title, keywords, abstracts, contents, and references, and then clean them to create the experimental dataset.
Then, the dataset $D$ is the article set from 51 journals, $D=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles. Each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{k},...,r_{K}\}$, and $K$ represents the length of the reference set.
3.2 Task Design
Since literature review writing primarily involves the collection and synthesis of relevant research, we design three tasks as follows to evaluate LLMs’ capabilities in different aspects of literature review writing.
  • Task 1: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
  • Task 2: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract $a_{i}^{g}$according to the research topic provided in the title and the keywords. The length of the generated abstract is to be the same as the length of the original abstract of the article. In this task, we evaluate whether LLMs can write an abstract based on the given topic.
  • Task 3: Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a short literature review $c_{i}^{g}$ according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ (with a total of $N$ citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
Three task prompts are shown in Appendix Table LABEL:at1.
3.3 Evaluation Metrics
Based on the type of generated text, we divide the evaluation of the model’s results into two parts: first, the hallucination rate of the references generated by LLMs, and second, a comparison of the generated context with human-written results, including two dimensions: factual consistency and semantic coverage.
Figure 1: Illustration of the experiment process.
Reference hallucination evaluation metrics. Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate true references. In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM. A higher accuracy metric indicates a lower hallucination rate.
For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task 1 and Task 3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the generated article title, authors, journal, year, volume, first page, and last page. Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
Next, we describe how to obtain $\{e_{d}\mid d=0,1,\dots,7\}$. First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines. Then, we obtain two candidate sets $Z_{t}$ and $Z_{a}$ respectively. Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $Z_{t}$ and $Z_{a}$. For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$. Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $z_{n}^{g}$ of $r_{n}^{g}$.
Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $z_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1). On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article. Here its title $T$ is correct or not, we give LLM a certain margin of error, as long as the title has a match rate of 80%, it will be judged as accurate, so we need another piece of information that is correct to ensure that the article is reliable. On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article. Therefore, when the other three elements in $r_{n}^{g}$ are matching, we also consider it as a reliable article.
$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}$ (1)
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(p_{i})$ across articles, as shown in Eq 3.
$Acc_{ref.}(p_{i})=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})$ (2)
$Acc.=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness. In  Agrawal et al. (2024)’s work, they determined by ground-truth labels assigned using the Bing search API. Encouraged by this work, we also calculate the title search scores of generated references for each LLM, formulated as follows.
$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}$ (4)
$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (5)
where $s_{p_{i}}$ is the number of the generated title that can be matched in the external scholar base. We search the generated title from an external academic search engine to estimate whether the title can be matched.
Context evaluation metrics. In this paper, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects. The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al. (2023). Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text. For example, we calculate the NLI score $Entail_{p_{i}}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
$Entail_{p_{i}}=\theta_{\text{NLI}}(a_{i}^{g},a_{i})=\begin{cases}1&\text{if }a_{i}^{g}\text{ entails }a_{i}\text{,}\\ 0&\text{otherwise}\end{cases}$ (6)
where $\theta_{NLI}$ denotes the NLI model. Finally, we obtain the NLI score $Entail$ for each model according to Eq 7.
$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}$ (7)
Additionally, we use commonly employed semantic similarity metrics and Key Point Recall (KPR) to calculate the semantic coverage between the context generated by LLMs and human-written context. Specifically, for Task 2, we apply cosine similarity and the ROUGE metric for semantic coverage evaluation. For Task 3, we use the ROUGE metric and KPR to measure the semantic coverage of the literature review generated by the LLMs relative to human-written content.
KPR, first proposed by  Qi et al. (2024), is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents. Since human-written literature reviews are lengthy and difficult to compare directly, we adopt the KPR method to measure the extent to which LLM-generated content covers the key points in human-written literature reviews. Specifically, we first use GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculate the coverage of these key points by the model-generated literature review as Eq. 8.
$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (8)
where $c_{i}^{g}$ denotes the literature review generated by LLMs.
Finally, we also concatenate key points and compute the ROUGE metric between the key points and $c_{i}^{g}$.
4 Experiments
4.1 Experimental Settings
Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1105 articles, the dataset distribution across journals as shown in Appendix B Figure 4.
LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3, and Llama-3.2-3B-Instruct.
In task 1 and task 3, we set $N$ as 10, each model generates 10 references. For the generated reference evaluation, we use Semantic Scholar as the external database. Recently, LLMs-as-judges has become more common Chen et al. (2024); Zheng et al. (2023); Shangyu et al. (2024). So, for Task 2, we employ TRUE Honovich et al. (2022), along with GPT-4o as NLI models for evaluating factual consistency in context; To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings. For task 3, we employ GPT-4o as the NLI model in Eq 8, and set $q$ as 10.
4.2 Main Results
We present the results for the three tasks in Table 12, and 3. The performance of different models on each task is analyzed as follows.
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5-Sonnet achieves the highest accuracy and title search rate, whereas Llama-3.2-3B exhibits the lowest performance on both metrics for Task 1. Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs. Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2. The graph shows that the majority of the correctly cited references for almost all LLMs involved in the task are concentrated between 2010 and 2020.
Results for Task 2. We use various metrics to evaluate model performance. As shown in Table 2, Claude-3.5-Sonnet consistently achieves the best overall performance across nearly all metrics. The average semantic similarity between the abstracts generated by Claude-3.5-Sonnet and human-written abstracts is 81.17%. Furthermore, Claude-3.5-Sonnet and DeepSeek-V3 exhibit the highest factual consistency, achieving a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.84%, respectively. Additionally, we observe that Llama-3.2-3B achieves the highest score on the ROUGE-L metric but does not exhibit a clear advantage on other metrics. This suggests that, given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
Results for Task 3. As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3. Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text (Gao et al., 2023). In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly. This suggests that the generated references and the literature review text impose constraints on each other.
On the other hand, we find that Claude-3.5-Sonnet achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5-Sonnet recalls the most claims from the human-written literature review. While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
  Models Task 1
$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
Qwen-2.5-72B 21.80 12.25 17.58
Llama-3.2-3B 16.62 3.45 6.95
DeepSeek-V3 56.04 46.33 50.66
GPT-4o 32.07 21.65 24.65
Claude-3.5-Sonnet 64.82 51.59 55.77
Table 1: The experimental results of the five LLMs in Task 1. “$S_{t}$” refers to the title search rate as defined in Eq 4, while “Acc.” represents the accuracy defined in Eq 3. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task 2
[…]
Qwen-2.5-72B 80.22 69.52 95.02 40.61 8.78 20.12
Llama-3.2-3B 79.28 62.39 92.14 40.35 8.96 20.52
DeepSeek-V3 80.96 78.55 96.84 41.13 8.98 20.33
GPT-4o 80.96 77.91 96.50 40.70 8.56 19.86
Claude-3.5-Sonnet 81.17 78.90 96.77 41.13 8.99 20.00
Table 2: Compare the performance of four LLMs on Task 2.
  Models Task 3
References Literature Review
[…]
Qwen-2.5-72B 40.02 28.91 33.64 38.82 29.95 9.01 15.14
Llama-3.2-3B 21.78 4.86 7.28 29.07 28.07 7.77 15.46
DeepSeek-V3 62.29 52.81 55.38 56.02 35.65 10.40 17.46
GPT-4o 60.05 50.62 54.16 59.18 30.78 9.72 15.54
Claude-3.5-Sonnet 66.43 59.06 63.06 62.32 28.59 8.90 14.41
Table 3: The experimental results of the four LLMs in Task 3. “$S_{t}$” refers to the title retrieval rate as defined in Eq 4.“Acc.” represents the accuracy defined in Eq 4. “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task1 Task3
Title Acc. Overlap Overlap(first author) Title Acc. Overlap Overlap(first author)
Qwen-2.5-72B 7.20 12.60 13.27 12.15 17.36 18.31
Llama-3.2-3B 5.97 8.48 8.67 8.01 8.28 8.44
DeepSeek-V3 14.28 19.72 20.50 20.44 26.79 27.30
GPT-4o 8.66 18.76 19.50 18.31 27.88 28.86
Claude-3.5-Sonnet 16.06 24.34 25.21 21.20 31.90 33.25
Table 4: Overlap rate of citations between LLM-generated and human-written references in Task 1 and Task 3. “Title Acc.” refers to the title overlap between LLM-generated and human-written references, while “overlap” represents the article overlap between LLM-generated and human-written references. “first author” refers to when evaluating the overlap of references, the author dimension only comparing the first author.
Figure 2: Distribution of LLM-generated true references over years.
4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5-Sonnet demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5-Sonnet, GPT-4o, and Qwen-2.5-72B follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. Accuracy for journal name, page, and author is also relatively high. However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions. In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions. However, overall, Llama-3.2-3B does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B, with particularly obvious gains in the author dimension. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension. Notably, the accuracy of DeepSeek-V3 and GPT-4o in certain dimensions approaches or even exceeds that of Claude-3.5-Sonnet. However, the performance of LLaMA-3.2-3B remains suboptimal.
4.4 Cross-Disciplinary Analysis
In this section, we compare the performance of LLMs across different disciplines. First, based on Dewey’s Decimal Classification, we categorize 51 journals into five disciplines: Biology, Chemistry, Mathematics, Physics, Social Science, and Technology. After categorization, there are 460 articles in the Biology category, 90 in Chemistry, 50 in Mathematics, 113 in Physics, 299 in Social Science, and 94 in Technology.
We then present bar charts in Figures  LABEL:f4, which illustrate the performance of different models across various tasks and disciplines.
First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline. Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065). These results indicate significant differences in performance across disciplines for each model except Llama-3.2-3B.
Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
Additionally, we examine the references accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c. It is evident that the accuracy of Claude-3.5-Sonnet, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5-72B and LLaMA-3.2-3B across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001), indicating significant differences in each models’ performance across the various disciplines.
Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d. The results from the figure indicate that the differences between models—Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B—are not significant across various disciplines, a finding that is also supported by statistical tests. Specifically, the p-values are as follows: Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).
4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5-Sonnet are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. We provide a more detailed comparison of the references across various dimensions in Appendix C.
4.6 Human Evaluation
To evaluate the reliability of our automatic assessment method for identifying hallucinated references, we conduct a comparative analysis involving 100 LLM-generated references. These references were assessed by three annotators and the final manual results were obtained by majority vote. The results demonstrated a kappa agreement of 0.71 between the automatic and human assessments, signifying a relatively high level of consistency and supporting the reliability of our method. Furthermore, when using human assessment results as the gold standard, the automatic assessment method achieved an accuracy of 86%, further validating its effectiveness.
5 Conclusion
In this paper, we present a framework to assess the literature review writing abilities of LLMs. This framework includes three tasks designed to evaluate LLMs’ literature review writing capabilities. The generated outputs are then evaluated from multiple dimensions using various tools, such as Semantic Scholar and NLI models, focusing on aspects like hallucination rate, semantic coverage, and factual consistency compared to human-written texts. Finally, we analyze the performance of LLMs in writing literature reviews from the perspective of different academic disciplines.
Limitations
In this paper, we evaluate the ability of LLMs to write literature reviews. However, instead of assessing from conventional perspectives, such as the fluency of the generated context or topic coverage, we compare the model-generated results with human-written ones. Additionally, when processing the LLM-generated results, we often need to handle abbreviations of author names and journal titles. While we have made efforts to address these details thoroughly, some discrepancies may still exist.
Appendix A Prompts for Tasks
Table 5: Prompts for tasks.
Prompt Content
Prompt 1 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to find the 10 studies that are most relevant to the research topic provided in the "Title" and the "Keywords" below.
Please cite the studies according to the following JSON format. There is no need to provide any explanation before or after the JSON output. Ensure that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.". { "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
Title: title
Keywords: keywords
Prompt 2 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write an abstract according to the research topic provided in the "Title" and the "Keywords" below. Please write the abstract for about xx words, according to the JSON format as follows. There is no need to provide any explanation before or after the JSON output. {"Abstract": ""}
Title: title
Keywords: keywords
Prompt 3 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write a literature review according to the research topic provided in the "Title", “Abstract” and "Keywords" below. The literature review should be about 1000 words long. I would like you to back up claims by citing previous studies (with a total of 10 citations in the literature review). The output should be in JSON format as follows: { "Literature Review": "xxx", "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
The "Literature Review" field should be about 1000 words. The "References" field is a list of 10 references, and ensures that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.".
Title: title
Keywords: keywords
Abstract: abstract
Figure 4: Statistics of dataset.
Appendix B Data Distribution
Statistics of the dataset are shown in Figure 4.
Appendix C Comparison of LLM-Cited and Human-cited References from Different Dimensions
We provide a more detailed comparison of the LLM-cited and human-cited references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5-Sonnet and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
Appendix D Discussion
We select four LLMs for task evaluation and find that Claude-3.5-Sonnet outperforms DeepSeek-V3, GPT-4o, Qwen-2.5-72B, and Llama-3.2-3B across all three tasks, particularly excelling in the task of generating accurate references. This advantage is likely influenced by the training data of each model. Additionally, we observed that each model has different strengths across disciplines. Overall, for the reference generation task, nearly all models perform better in Mathematics, while their performance is weaker in Chemistry and Technology. However, when writing abstracts, all models exhibit the lowest factual consistency in Social Science, as indicated by the entailment scores, compared to human-written texts.
When comparing the references generated by the models in Task 1 and Task 3, we find that in Task 3, nearly all models generate more accurate references. This suggests that LLMs cite references during the writing process, which improves the authenticity of the references. Moreover, the inclusion of the first author’s name in the generated context also enhances the accuracy of the author dimension.
In the future, we will deploy this evaluation framework on Hugging Face to enable real-time assessments of newly released LLMs.